OpenStudy (anonymous):

In this problem Write two functions, calledcountSubStringMatch and countSubStringMatchRecursive that take two arguments, a key string and a target string. These functions iteratively and recursively count the number of instances of the key in the target string. what would be the iterative solution?

OpenStudy (anonymous):

I thought of this solution and i think it is of the category recursive from string import* def countrecursive(target,key): y=0 x=0 z=0 while(y!=-1): y=find(target,key,x) if(y>=0): x=y+1 z=z+1 return z

OpenStudy (anonymous):

That's not recursive. That is iterative. A recursive function calls itself, while an interative one uses a loop instead.

OpenStudy (anonymous):

ohh thanks...I will try again